OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+3=-2x+8 5x=5 x=1

OpenStudy (anonymous):

@melanie083194 medal please